UT Dec 04 2018

D-Night + MODS AGW Transform

OSA: D.G. Huerta

ISA: D. Thompson (Tucson)

ISp: E. Solheid

Summary:

Barry: HD 216807, NGC 7727 HK & zJ

HD 18571, NGC1110 HK & zJ

SDSSJ0228-0901 zJ only

HD18581 (really HD18571?) zJ only

Mark: G191b2b, V392Per (1x @ new PA)

V959Mon orig guide* was double,

V959Mon @2nd PA because 1st was 40deg off

<did not get second standard>

MODS transform

MODS1 the on-axis star stayed pretty much in the same spot

MODS2 the on-axis star was all over the place, no idea why

MODS Cals

grpixflats.cal done

grpixflats_b.cal done

grpixflats_r.cal error setting nimgs?

LUCI Cals

Bino_LUCI_Cals_HKzJ_2pt0slit.xml done

Bino_LUCI_Cals_0pt75slit.xml has incorrect mask!

Cals_Bino_G210_0pt5_2pt27.xml done

Oddities:

Failed co-pointing check on 75" DETXY sync offset?

LUCI Telluric scripts changed crossed filters for final acq image (OT)

LUCI2 AFC spots seem in poor focus ... recheck CamGrat focus values ASAP

Script/OT problem? OT showed HKspec/1.93um but script had zJspec instead (IT#7499?)

modsDisp won't run if you are in ~/Desktop ?

Errors during MODS grpixflats_r.cal and diBias scripts

>> Check that necessary calibrations have been taken!

Transform check: MODS2 the on-axis star was all over the place!

Other Notes:

All programs should be provided with the OT file...difficult to fix simple mistakes without!


Details:

cd /home/lbto/Service_Observing/DTime_UTDec04_2018/

evince Visibilities_Dec03_Targets.pdf

Pretty clear @ sunset. Zodi is easily visible on all-sky cam data once dark

thin clouds all around @ sunrise

### Sunset-Twilight LUCI-2 AO checkout with updated guidelines

LUCI subwindow:

subwin clear

subwin sw 1 xs ys xsize ysize

subwin auto on

subwin off

subwin clear

UT 01:22 Completed (no LUCI data taken)

|| Note: If LUCI images are not likely to be obtained, don't bother moving the

|| N30_FieldStop mask into the FPU, unless AO observations will immediately follow.

### Start of Night - NGC 7727 - Binocular LUCI LS (0.5" G210K)

/home/observer/Service_Observing/DTime_UTDec04_2018/Rothberg/CO_Merger/

README: BR_Merger.README

Notes: Observe the telluric first before the science script.

### HD 216807 22 56 05.49 -12 08 47.4 V=9.97 Telluric A0V

Script: Telluric_for_NGC7727_Bino.xml

01:34 Script sent

>> Telluric script changed crossed filters in the middle of the acquisition!

01:45 acq done, starting spectra

1.15" on guiders, 0.99 DIMM

01:49 Script done

files up to 0005 / 0011

AFC spots on LUCI2 image look less distinct than the ones on LUCI1.

Need to re-check CamGrat focus offsets!

### NGC 7727 23 39 53 -12 17 35 1855 2.307 90.0

01:49 Script sent: NGC7727_Bino.xml

eog NGC7727_K_Finder.jpeg

Files 0006-0009 0012-0015 acquisition

did not need sky frame, core of galaxy more than bright enough

~11k ADU/DIT

L2 subtraction took a long time (stopped reset and restarted after 2 min)

-probably lost 5min total on this

FieldStop is offset on LUCI2

Corrected after acq and before starting spectra using 0012

Error - co-pointing limit

Preset still active

LUCI Telsvc sync offset to 0.0,0.0 DETXY sent to return to acquired pos'n

co-pointing check turned off

reset queue on LUCI, skip to spectrum and go

...good thing John was here, I might have re-acquired

02:13 spectra started

end of first spectrum: GCS shows 0.95 L1 & 1.08 L2

LUCI2 AFC spots are fairly out of focus w/tails (top and bottom of slit)

files start at 0010 L1 and 0016 L2

03:18 Done. Last files 0015 0021

### NGC 1110 - Binocular LUCI LS (2" G200 HK then zJ)

/home/observer/Service_Observing/DTime_UTDec04_2018/Rothberg/IMBH/

Observe the telluric first before science target.

HD 18571 02 59 16 +01 14 40 V=8.63 2" slit Telluric for SDSS J0228-0901 and NGC1110

NGC 1110 02 49 09 -07 50 23 20.6 0.0044 2"

03:18 sent script HD18571_Telluric_HKzJ_Bino.xml

Script uses OH_1060+z for acq imaging, HeI+J for confirming image

-0.25" dx on L2 from telsvc

Barry fixed a problem from the OT where 1.93um but zJspec was in config

reload, skip to and go...

Spectra 0019-0020 0025-0026 HKspec

0021-0022 0027-0028 zJspec

03:41 done

03:42 sent script NGC1110_HKzJ_Bino.xml

03:53 network outage

04:06 science spectra started

05:18 Network dropout again

05:51 Done

last spectra 0050 0056

05:51 sent script SDSSJ0228-0901_zJ_Bino.xml

failed preset - pointing correction needed

05:56 sent script again

FieldStop is offset on both LUCIs

Corrected after acq and before starting spectra

06:18 Started spectroscopic section

06:54 done last spectra 0061 0067

### Another visit to std, just zJspec

06:54 sent script HD18581_Telluric_zJ_Bino.xml ??

does HD18571, not 18581 ??

07:09 Done with LUCI ...

FC off, blind+blind, mirror, N375 both sides.

DGH reconfiguring

##########################################################

### MODS Startup

modsDisp is opening 4 DS9s and exiting with an error:

Starting the modsDisp session...

Launching MODS1Blue ds9 display, please wait...

Launching MODS1Red ds9 display, please wait...

Configuring the MODS1Blue ds9 display...

Traceback (most recent call last):

File "/home/MODSeng/bin/modsDisp", line 346, in <module>

m1bDisp.set('height 800')

File "/lbt/mods_runtime/python-mods/lib/python2.6/site-packages/ds9.py", line 368, in set

return xpa.xpaset(self.id, paramlist, s, blen, 1)

File "/lbt/mods_runtime/python-mods/lib/python2.6/site-packages/xpa.py", line 149, in xpaset

if errmsg: raise ValueError, errmsg

ValueError: XPA$ERROR can not find channel named "sock12" (DS9:MODS1Blue c0a83a21:49627)

>> email sent to OK+SciOps about this

>> Barry's suggestion "fixed" it...modsDisp won't run from

~/Desktop, but will from ~/

UT 01:30 Both MODS have been awakened

07:14 Both MODS set to Observing mode

homed both AGw

07:38 acqBinoMODS G191b2b.acq

7:42 modsAlign -r mods1r.20181204.0001.fits

7:44 modsAlign -r mods2r.20181204.0001.fits

07:47 execBinoMODS G191b2b.obs

PA=0 PAR=170.7

Good seeing...0.52" on guiders

Did 3 exposures in each mode: Dual grating, Red grating, Blue grating

--> NO prism data!

### V392Per

evince v392per_PA_63deg.pdf &

8:23 acqBinoMODS v392per_new.acq

8:29 modsAlign -y 11 mods1r.20181204.0009.fits mods1r.20181204.0010.fits

8:30 modsAlign -y 9 mods2r.20181204.0009.fits mods2r.20181204.0010.fits

8:33 execBinoMODS v392per.obs

looks like both broad and narrow Halpha emission

PAR 117 @ end of script, decreasing

08:55 Network dropout again

### V959Mon

evince v959mon_PA_-30deg.pdf &

8:57 acqBinoMODS v959mon.acq

guide star is a close double

9:05 modsAlign -y 11 mods1r.20181204.0015.fits mods1r.20181204.0016.fits

9:06 modsAlign -y 9 mods2r.20181204.0015.fits mods2r.20181204.0016.fits

>> started .obs before taking the confirming exposure!

9:06 execBinoMODS v959mon.obs

PAR=1.0 @ start, increasing

looks like a shell around this one.

PAR=17 @ end

>> OT file was not provided ... this makes it hard to respond to both guide star

>> problems and updates to the PARANGLE based on time of observation!

Collimation seems to be holding OK

...maybe not. Was fuzzy @ end, but did not seem to run away

### Run again, better PA and new guide*

09:45 acqBinoMODS v959mon_PA33.acq

9:50 modsAlign -y 11 mods1r.20181204.0020.fits mods1r.20181204.0021.fits

9:51 modsAlign -y 9 mods2r.20181204.0020.fits mods2r.20181204.0021.fits

manually adjusted MODS1 -0.15"X and MODS2 +0.25"X and absorbed

9:55 execBinoMODS v959mon.obs

PAR=21.5 @start

PAR=33.9 @end

09:59 another network dropout

### MODS transform collection

UT 10:35

zeropoints are already updated in PCS

Olga will update the AGW.ini file with the new trasform and restart the services.

>> both already done before the start of the night, per emails

to configure MODS

cd /home/modseng/modsScripts/Support

execBinoMODS transformInit_PointingMask.pro

>> done

>>> Shouldn't this be red-red now that MODS2R is fixed?

Preset to field, in one IDL window:

iif_register, /TCS

iif_preset, Stone_E, MODE=Active, GS=50, TS=50, PA=180, /RIGHT

iif_preset, Stone_E, MODE=Active, GS=50, TS=50, PA=180, /LEFT

iif_binoc_preset

Once you are collimated and happy, in the MODS GUIs, in the command line part,

in MODS guis command lines: agw reset

iif_register, /LEFT

iif_register, /RIGHT

transform_collect: https://wiki.lbto.org/Commissioning/ComTools#Guide_Probe_Transformation

10:55 Its working now, both sides.

In MODS1 the same star stayed pretty much in the same spot in all images

In MODS2 the on-axis star was all over the place. No idea why.

12:47 done

12:50 Start it over for at least a few *s

MODS1 looks OK, MODS2 is way off.

Field at EL=32! Part of the problem?

06:04 MST Done...parking telescope

### Calibrations

execBinoMODS grpixflats.cal

MODS1 vacuum pressure on blue channel went red but recovered

execBinoMODS grpixflats_b.cal

execBinoMODS grpixflats_r.cal

** Error: nimgs Cannot change this parameter during image acquisition

** Abort, Retry, or Ignore? >

source doBias

06:15 MST network dropout again

06:28 Another network dropout

Lengthy. Seemes like the local computer was also frozen (rm507-3)

doBias (just 8k and 3k)

MODS2 hung during readout of blue channel?

** ERROR: Command 'blue abort' timed out after 30 seconds

... abort sequence terminated with errors.

** Binocular MODS2 script aborting on fatal error.

### LUCI1

>> LUCI1 rotator was at 288.7 ... set to 341.0

06:35 MST Bino_LUCI_Cals_HKzJ_2pt0slit.xml

Files: 0066-0099 L1 ; 0072-0105 L2

07:05 MST Bino_LUCI_Cals_0pt75slit.xml

-- ABORT! (did not change mask ??)

>> Script has incorrect mask in it !!

07:12 MST Cals_Bino_G210_0pt5_2pt27.xml

Files: 0100-0120 L1 ; 0111-0131 L2

-- BarryRothberg - 04 Dec 2018
Topic revision: r1 - 04 Dec 2018, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback